Xlog2+log5=1 find x
THANK YOU

Answer by stanbon(75887)

Xlog2+log5=1 find x
-----
log(2^x) + log(5) = 1
-----
log[5*2^x] = 1
-----
5*2^x = 10
---
2^x = 2
--
x = 1
------------
Cheers,
Stan H.